Old World Charm, Inc. specializes in selling scented candles. The company has established a policy of reordering inventory every other month (which is 6 times per year). A recently employed MBA has considered New England's inventory problem from the EOQ model viewpoint. If the following constitute the relevant data, what is the extra total cost of the current policy compared with the total cost of the optimal policy? Enter your answer rounded to two decimal places. Do not enter $ or comma in the answer box. For example, if your answer is $12,300.456 then enter as 12300.46 in the answer box.

Ordering cost = $10 per order

Carrying cost = 20% of purchase price

Purchase price = $15 per unit

Total sales for year = 1,000 units

Safety stock = 0
Answer #1

SOLUTION:-

Annual Quantity Demanded = 1000 units

Ordering Cost per order = $10 per order

Holding Cost per unit per year = 20% of Purchase Price = 20% of $15 = $3 per unit per year

EOQ - Cost under Optimal Policy:

EOQ = Square root of (2 x Annual Quantity Demanded x Ordering Cost per order) / (Holding Cost per unit per year)

= Squareroot of (2*1000*10)/3

= Squareroot of 20,000/3

= Squareroot of 6,666.67

= 81.64966 units

No. of orders to be made in a year = 1,000/81.64966

= 12.247 orders

= 12.247 orders

Annual Ordering Cost = 12.247*10

= $122.47

Annual Carrying Cost = (81.65/2)*3

= $122.47

Total Inventory Cost if we order in EOQ = Annual Ordering Cost + Annual Carrying Cost

= $122.47+ $122.47

= $244.94

Current Policy - Cost under Existing Policy:

No. of orders in a year = 6

Oder Quantity per order = 1,000/6 = 166.66 units

Annual Ordering Cost = 6*10

= $60

Annual Carrying Cost = (166.66/2)*3

= $250

Total Inventory Cost if we order in EOQ = Annual Ordering Cost + Annual Carrying Cost

= $60 + $250

= $310

.The extra total cost of the current policy compared with the total cost of the optimal policy = $310- $244.94

= $65.06